OpenStudy (anonymous):

Is the ordered pair (0,5) a solution of the system of linear equations: x + y = 5 -5x + 2y = 10 a. yes b. no Im really not sure on this

OpenStudy (anonymous):

Do you know how to solve systems?

OpenStudy (anonymous):

I would be lying if i said i did lol my school doesn't explain anything :(

OpenStudy (anonymous):

First take the first equation, and solve it for y. y = 5-x Then we plug this in to the next equation and solve for the value of y. -5x + 2(5-x) = 10 -5x + 10 - 2x = 10 -7x + 10 = 10 -10 -10 -7x = 0 /-7 /-7 x = 0 Now we plug x = 0 back in to the first equation: 0+y = 5 y = 5 Ordered pair: (0,5) SO yes it is a solution. Does this make sense?

OpenStudy (anonymous):

Please let me know if it does not.

OpenStudy (anonymous):

OH thank you!!! that made perfect sense :)

OpenStudy (anonymous):

Awesome tag me for any more help.

OpenStudy (anonymous):

Okay thank you